OpenStudy (anonymous):

HELPPPPPPPPPPPPPPP! A company uses two vans to transport workers from a free parking lot to the workplace between 7:00 and 9:00 a.m one van has 9 more seats than the other the smaller van makes two trips every morning while the larger one makes only one trip the two vans can transport 69 people maximum how many seats dose the larger van have let x be the seats in the small van and y be the seats in the large van. how many seats does the larger van have?

OpenStudy (anonymous):

@SolomonZelman @AriPotta @dejmccray @Ghostgate @freckles

OpenStudy (anonymous):

@Maddox1000 @daphne101 Are you guys here?

OpenStudy (anonymous):

dont worry imm thinking

OpenStudy (anonymous):

thankyou

OpenStudy (ghostgate):

I would think it would be represented like this: 2x + y + 9 = 69

OpenStudy (ghostgate):

Then of course you simplify and figure out the answer.

OpenStudy (anonymous):

@Ghostgate is correct

OpenStudy (anonymous):

i need help with simplifying it @Ghostgate

OpenStudy (anonymous):

this is how i do it

OpenStudy (ghostgate):

First you take the equation: 2x + y + 9 = 69 Subtract 9 from both sides which cancels out 9. So what is 69 - 9?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

first you said that the small bus goes 2 times well then figure out if that one has the 9 seats or the large one

OpenStudy (ghostgate):

Now your equation is: 2x + y = 60 Divide both sides by 2 which cancels out the 2 and you're left with x. So what is 60 divided by 2.

OpenStudy (anonymous):

then u can divide your answer to @Ghostgate by 2 for 2 large buses

OpenStudy (anonymous):

30

OpenStudy (anonymous):

so that would be the answer!

OpenStudy (anonymous):

youre good @Ghostgate

OpenStudy (ghostgate):

That would sound right, but most commonly you need either x or y on one side and the number on the other. Like: x = 30, but since we don't have that, I am now puzzled.

OpenStudy (anonymous):

its alright i put 30 and it was wrong but its alright i aprreciate the help alot! if i have another question can i @ you guys @Ghostgate @daphne101

OpenStudy (ghostgate):

Wait, I think 'y' is acting as a tricky x. Is the answer x = 10?

OpenStudy (anonymous):

i already anserwed i cant go back @Ghostgate

OpenStudy (ghostgate):

Oh, well I believe it was acting as a x. Why? Because as it listed the larger van(y) only had a difference of having 9 more seats compared to the other, but the smaller van made 2 trips making it 2x, which separated them. So, when removing the 9, I wasn't thinking of it being an extra x. In other words: 3x = 60, divide by 3 and you get 20. And when you plug that in: 2(20) + y + 9 = 69 40 + y + 9 = 69 49 + y = 69 y = 20 Now plug it in: 2(20) + 20 + 9 = 69 40 + 29 = 69 69 = 69

OpenStudy (anonymous):

thankyou i might need this agin

OpenStudy (ghostgate):

You're welcome! ;) - I'm glad I could help!